Consider a market with two risk-free zero-coupon bonds, A and B. Their respective maturities are 1 and 2 years, and their market prices are 97.0874 and 95.1814 (expressed as percentage of the face value).

(A) Calculate the implied forward rate between years one and two, f1;2 .

(B) According to the Pure-Expectations Theory, what are the investors views about the one-year rate r1 one year from now?
Answer #1
A) Implied forward rate = 97.0874/95.1814-1 = 2.00%
B) Current 1 Year rate = 100/97.08-1 = 3.01%
Current 2 Year rate = (100/95.1814)^(1/2)-1 = 2.50%
1 Year rate 1 Year from now = 1.025^2/1.0301-1 = 1.99%